OpenStudy (anonymous):

-5x+12y=20, x-2y=-10 solve by using either the elimination method or the substitution method.

OpenStudy (anonymous):

-5x+12y=20, x-2y=-10 solve by using either the elimination method or the substitution method.

OpenStudy (anonymous):

-5x+12y=20, x-2y=-6 solve by using either the elimination method or the substitution method.